A 70 kg Person sitting on a frictionless stool Is Please explain the process.
Answer #1

The initial angular momentum is equal to final angular momentum so that angular momentum is conserved.

I_1 \omega_1 =I_2\omega_2

(\frac{1}{2}mr^2 + 2mr_1^2) \omega_1 =(\frac{1}{2}mr^2 + 2mr_2^2)\omega_2

(\frac{1}{2}(70)(0.2)^2 + 2(70)(0.8)^2) (3) =(\frac{1}{2}(70)(0.2)^2 + 2(70)(0.2)^2)\omega_2

\omega_2= 39 rad/s

b)

The change in rotation kinetic energy is

\Delta K = \frac{1}{2}I_2\omega_f^2 - \frac{1}{2}I_1\omega_i^2

\Delta KE = \frac{1}{2}(\frac{1}{2}(70)(0.2)^2 + 2(70)(0.2)^2)(39)^2 -\frac{1}{2} (\frac{1}{2}(70)(0.2)^2 + 2(70)(0.8)^2) (3)^2

\Delta KE = 4914J